OpenStudy (narissa):

what property is demonstrated a(3*5*7=7*5*3 b)(3+5)+7=3+(5+7)

OpenStudy (anonymous):

\[3\times 5\times 7=7\times 5\times 3\] numbers changed order that is called the "commutative property" of multiplication

OpenStudy (anonymous):

for the second one, the numbers stay in the same order, but the parentheses move that is the "associative property"

OpenStudy (narissa):

thanks

OpenStudy (anonymous):

yw